A solid spherical ball of mass 1.72 kg and radius 0.16 m rolls along a smooth, level surface with a speed of 2 m/s.  What is the ball
Answer #1

moment of Inertia of sphere, I = (2/5)*m*R^2

= 0.4*1.72*0.16^2

= 0.0176 kg.m^2

angular speed, w = V/r

= 2/0.16

= 12.5 rad/s

Rotational KE = 0.5*I*w^2

= 0.5*0.0176*12.5^2

= 1.375 J